a seaplane flying with the wind flew from an ocean port to a lake,a distance of 240 mi in 2h.Flying against the wind,it made the trip from the lake to the ocean port in 3h.Find the rate of the plane in calm air and the rate of the wind.
Answer by TimothyLamb(4379)

s = d / t
---
a + w = 240/2 = 120
a + w = 120
a = 120 - w
a - w = 240/3 = 80
---
a - w = 80
120 - w - w = 80
40 = 2w
---
w = 20
a = 100
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php